OpenStudy (anonymous):

in 1990 retail sales at bookstores were about 7. 4 billion. in 1997 retail sales at bookstores were about 11.8 billion. write a linear model for retail sales s (in billions of dollars) at bookstores from 1990 through 1997. let T represent the number of years since 1990. Then estimate the retail sales at bookstores in 2012.

OpenStudy (math&ing001):

You're given 2 points (1990 , 7.4) (1997,11.8) and you're asked to find a linear model (equation of a line from them. First you find the slope :\[m=\frac{ changein y }{ changein x }=\frac{ 11.8-7.4 }{ 1997-1990 }=\frac{ 4.4 }{ 7 }\] Then the equation would be y-y1=m(x-x1) ie. y-7.4 = 4.4/7 (x-1990) So f(x) = 4.4/7 (x-1990) + 7.4 Simplify it and it should give you the linear model you're looking for.